Use the function h(t)=-16t^2+v0t+s
A projectile is launched from a building 256 feet tall. If the initial velocity of the projectile was 96 feet per second, how long will it take the projectile to land on the ground?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Use the function h(t)=-16t^2+v0t+s
A projectile is launched from a building 256 feet tall. If the initial velocity of the projectile was 96 feet per second, how long will it take the projectile to land on the ground?
------------
Impact at h(t) = 0
Solve for t.

Answer by stanbon(75887)

Use the function h(t)=-16t^2+v0t+s
A projectile is launched from a building 256 feet tall. If the initial velocity of the projectile was 96 feet per second, how long will it take the projectile to land on the ground?
-----
Equation:
0 = -16t^2+96t+256
-----
Divide thru by -16 to get:
t^2 - 6t - 16 = 0
(t-8)(t+2) = 0
Positive solution::
time = 8 seconds
--------------
Cheers,
Stan H.
-------------